# Quantifying the Impact of Incompleteness on Identifying and Interpreting Galaxy Protocluster Populations with the TNG-Cluster Simulation

### v1

@@ Appendix A Observational Feasibility of Protocluster Definition Employed in This Work
In Fig. 14, we explore how our definition of simulated protoclusters – galaxies that will reside within the virial radius ($R_{\rm{200c}}$) at $z=0$ – compares with observational scenarios, where such information is inaccessible. Specifically, we compare the contamination fraction, defined as the ratio of interlopers to the sum of protocluster members and interlopers, as a function of separation from the center of mass of the protocluster. Here, interlopers are defined as galaxies that reside within the Lagrangian volume (traced by the protocluster members) but will not reside within $R_{\rm{200c}}$ at $z=0$. In contrast, protocluster members are galaxies that will reside within $R_{\rm{200c}}$ at $z=0$. Both interlopers and protocluster members are constrained to have stellar masses greater than $10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$.
@@
Appendix B Table of Spectroscopically-confirmed Protoclusters
Table 1: Spectroscopically-confirmed Protoclusters at $z>2$ with $N_{\mathrm{spec}}>10$
@@ Appendix B Table of Spectroscopically-confirmed Protoclusters
c Number of spectroscopically-confirmed protocluster members.

### v2

@@ Appendix A Observational Feasibility of Protocluster Definition Employed in This Work
In Fig. 14, we explore how our definition of simulated protoclusters – galaxies that are progenitors of those that reside within the virial radius ($R_{\rm{200c}}$) at $z=0$ – compares with observational scenarios, where such information is inaccessible. Specifically, we compare the contamination fraction, defined as the ratio of interlopers to the sum of protocluster members and interlopers, as a function of separation from the center of mass of the protocluster. Here, interlopers are defined as galaxies that reside within the Lagrangian volume (traced by the protocluster members) but will not reside within $R_{\rm{200c}}$ at $z=0$. In contrast, protocluster members are galaxies that will reside within $R_{\rm{200c}}$ at $z=0$. Both interlopers and protocluster members are constrained to have stellar masses greater than $10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$.
@@
Appendix B Table of spectroscopically confirmed Protoclusters
Table 1: spectroscopically confirmed Protoclusters at $z>2$ with $N_{\mathrm{spec}}>10$
@@ Appendix B Table of spectroscopically confirmed Protoclusters
c Number of spectroscopically confirmed protocluster members.